OpenStudy (anonymous):

derivative of lnx/x^5

OpenStudy (anonymous):

you have formula for this. just apply

OpenStudy (anonymous):

d/dx lnx is 1/x

OpenStudy (anonymous):

but what do I do with the x^5 in the denominator?

OpenStudy (anonymous):

[f(x)/g(x)]' = \[\frac{ g(x)*f(x)'-f(x)*g(x)' }{ [g(x)]^2 }\]

OpenStudy (anonymous):

your f(x) =ln x; your g(x) =x^5

OpenStudy (anonymous):

ah, I thought I could maneuver it without the quotient rule, I guess not. Thanks

OpenStudy (anonymous):

ok. take your time.